if secA-tanA=a then sinA=
options
  1. 1-a2/1+a2
  2. 1+a2/1-a2
  3. 2a/1-a2
  4. 2a/1+a2
  5. 1+a2/1-a2

Grade:12th Pass

1 Answers

Arun
6 years ago
we know that sec^2 A - tan^2 A =1 and it is given that 
sec A -tan A =a

dividing equation 1 with 2, we get 

sec A + tan A = 1/a
also sec A - tan A = a
 adding and subtracting these equations, we get 

sec A = (a + 1/a )/2 
& tan A = (1/a -a)/2 

thus, sin A = (sin A/ cos A) / (1/cos A) = tan A / sec A =( 1-a^2) / (1+ a^2)